Find:Differential to approximate value of cube root of 3.3

1
Expert's answer
2021-11-23T16:55:37-0500

Let "f(x)=x^{\frac 1 3}" . Formula for approximate calculation using differential is presented below

"f(x_0+Δx)\approx f(x_0)+df(x_0)" , where "df(x_0)=f'(x_0)*\u0394x"

"f'(x)={\frac 1 3}x^{-{\frac {2} 3}} = {\frac 1 3}*{\frac 1 {(x^2)^{\frac 1 3}}}"

Now we have to present number 3.3 as "x_0+\u0394x" such that "f(x_0)" is a perfect cube and "\u0394x" is relativately small. We can present it as 3.3 = 1 + 2.3, but 2.3 is a relativately big number, so the approximation error will be significant. Since "3.3^{\frac 1 3}={\frac {33^{\frac 1 3}} {10^{\frac 1 3}}}" , then we can present 3.3 as "{\frac {33} {10}}"

and calculate both f(33) and f(10) using differentiation

"f(33)=f(27+6)=f(27)+f'(27)*6=27^{\frac 1 3}+{\frac 1 3}*{\frac 1 {(27^2)^{\frac 1 3}}}*6=3+2*{\frac 1 9}={\frac {29} 9}"

"f(10)=f(8+2)=f(8)+f'(8)*2=8^{\frac 1 3}+{\frac 1 3}*{\frac 1 {(8^2)^{\frac 1 3}}}*2=2+2*{\frac 1 3}*{\frac 1 4}=2+{\frac 1 6}={\frac {13} 6}"

"f(3.3)={\frac {f(33)} {f(10)}}={\frac {29*6} {9*13}}=1{\frac {19} {39}}\approx1.49"

Lets check the accuracy

"1.49^3=3.307959" . The obtained number is close to 3.3, so we answer is very accurate